OpenStudy (anonymous):

matthhew loaned three friends a total of $27. all he can remember is that he loaned the second friend $5 more than the first and the third friend twice as much as the second. write and solve an equation to detemine how much he loaned each friend.

OpenStudy (anonymous):

27$ = first + second + third second = first + 5$ third = 2 * second = 2 * (first + 5$) 27$ = first + (first + 5$) + 2*(first + 5$) 27$ = 4 first + 15 12$ = 4 first 3$ = first second = 3$ + 5$ = 8$ third = 2 * 8$ = 16$ check again: 27$ = 3$ + 8$ + 16$ looks good to me ;)

OpenStudy (anonymous):

how did you get 3 16 8 and 15

OpenStudy (anonymous):

{3, 8, 16} is the solution. The 15 in the second row, is cause of: 27$ = first + (first + 5$) + 2*(first + 5$) 27$ = first + first + 5$ + 2*first + 2*5$ 27$ = first + first + 2*first + 5$ + 10$ 27$ = 4 first + 15 ok?

OpenStudy (anonymous):

does 2x repersent the second friend

OpenStudy (anonymous):

no the second friend got 5$ more than the first - so [first+5$] the third got double the amount the second got, so 2*[first+5$] all in all, he loaned first + second + third....

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

you're welcome :)